OpenStudy (cookiemonster18):

1.Solve the system of equations algebraically.Show all your steps. y=x^2+2x y=3x+20

OpenStudy (freckles):

x^2+2x=3x+20 is a quadratic equation write in standard form ax^2+bx+c=0 you can try to factor ax^2+bx+c or just use quadratic formula

OpenStudy (freckles):

there is also completing the square if you wish we just put it in the form x^2+bx=k and add (b/2)^2 on both sides giving x^2+bx+(b/2)^2=k+(b/2)^2 then this let's us write that one side as a square (x+b/2)^2=k+(b/2)^2 then so on...

OpenStudy (cookiemonster18):

okay thank you @freckles do you think you can lead me threw if by telling me to plug them all in and tell me if I got it right or not and help a little bit more. I don't know how to solve this correctly so I want to make sure I get it right when I do

OpenStudy (freckles):

ok please let me know what you after putting it in standard form

OpenStudy (freckles):

like what would be the first step in moving everything to one side?

OpenStudy (freckles):

you have two steps to perform to do do this

OpenStudy (freckles):

either one can go first

OpenStudy (freckles):

are you still there?

OpenStudy (cookiemonster18):

ax^2+bx+c=0 is the standard form? and I plug 3x^+20+?=0???

OpenStudy (cookiemonster18):

I'm sorry I honestly really don't have no idea how to solve or do thi . This why I need help learning how

OpenStudy (freckles):

\[x^2+2x=3x+20 \] we are trying to get everything to one side your first could have been to subtract 3x on both sides \[x^2+2x \color{red}{-3x}=3x+20\color{red}{-3x}\] redordering so like terms are together: \[x^2+2x-3x=3x-3x+20\] \[x^2-x=0+20\] \[x^2-x=20\] You still like one more step to have everything on one side together

OpenStudy (cookiemonster18):

What and where is the last step? Question does that solve one of the equations for the problem or both? I'm kinda confused now

OpenStudy (cookiemonster18):

@freckles ....you there???

OpenStudy (freckles):

have you moved everything to one side yet?

OpenStudy (freckles):

there is one more step to do so

OpenStudy (freckles):

do you see that 20 on the side what do we need to do to both sides so that one side is 0